Using Standard Normal Distribution, find the posibilitity

1. P( 0 < Z < 2.16 )
2. P( -1.87 < Z < 0)
3. P( -1.63 < Z < 2.17)
4. P( 1.72 < Z < 1.98)
5. P( Z > 1.77)
6. P( Z > - 1.02)
7. P (Z < 2.03)
Answer #1


1. P( 0 < Z < 2.16 )

= P(z< 2.16) - P(z< 0)
= 0.9846 - 0.5

= 0.4846


2)

P( -1.87 < Z < 0)

= P(z< 0) - P(z< -1.87)
= 0.5 - 0.0307
= 0.4693

3)

. P( -1.63 < Z < 2.17)
= P(z< 2.17) - P(z< -1.63)
= 0.9850 - 0.0516

= 0.9334

4)

P( 1.72 < Z < 1.98)
= P(z< 1.98) - P(z< 1.72)
= 0.9761- 0.9573
= 0.0189


5)

P(z> 1.77)
= 1 - P(z < 1.77)
= 1 -0.9616
= 0.0384

6)

P(z> -1.02)
= 1 - P(z< -1.02)
= 1 - 0.1539
= 0.8461

70

P(z< 2.03)
= 0.9788